zarkam21:

If cost = -0.42 and cosw = -0.80, both t and w are positive, and the angles determined by t and w are in quadrant 3, then which of the following statements is true? (2 points) a. t > w b. w > t c. Cannot be determined

zarkam21:

@Angle

Angle:

sorry for the messy drawing 

zarkam21:

A

Angle:

yup

zarkam21:

If sint = 0.29 and sinw = 0.43, both t and w are positive, and the angles determined by t and w are in quadrant 2, then which of the following statements is true? (2 points) a. t > w b. w > t c. Cannot be determined

zarkam21:

B?

Angle:

yes :)

zarkam21:

Convert 3.75 radians into degrees to the nearest two decimal places 214.86 degrees

Angle:

correct ^_^

zarkam21:

Convert 153 into radians to the nearest two decimal places. 2.67 radians

Angle:

awesome :D

zarkam21:

Convert 64.32 into degrees, minutes, and seconds.

1 attachment
zarkam21:

3685.26 degrees 221115.87 minutes 1.3267E+7 Seconds

Angle:

I think they mean 64 degrees + ___ minutes + ___ seconds like you would tell time: it is 0 hours, 59 minutes, and 30 seconds right now

zarkam21:

oh okay so it would be 64 degrees 59 minutes 30 seconds

Angle:

ehhh I was giving an example x'D it's not actually 59 minutes, 30 seconds we have to turn 0.32 degrees into ___ minutes + ___ seconds

zarkam21:

64 degrees 19 minutes 12 seconds

Angle:

perfect :)

Angle:

>I don't know why wolframalpha says 64 degrees, 19 minutes, 11 seconds >.>

zarkam21:

No idea, i'm just using google conversions

zarkam21:

1 attachment
zarkam21:

since 1 minute = 1/60 degree and one second = 1/60 minute = 1/3600 degree

Angle:

yup